# KG-TRACES: Enhancing Large Language Models with Knowledge Graph-constrained Trajectory Reasoning and Attribution Supervision

### v1

@@ 4.1 Experimental Setups
We conduct comprehensive experiments to evaluate the effectiveness and generalization ability of KG-TRACES across both general-domain and domain-specialized (medicine) reasoning tasks. We evaluate the general reasoning ability of KG-TRACES on two open-domain multi-hop KGQA benchmarks: WebQuestionsSP (WebQSP) [41] and Complex WebQuestions (CWQ) [29]. To evaluate cross-domain generalization, we additionally use GenMedGPT-5k [15], a medical QA benchmark constructed from ChatGPT-patient interactions and supported by a curated medical KG of symptoms, diseases, drugs, and treatments [34]. The details of the datasets are provided in Appendix A.
@@ B.2 Inference Prompts
These inference prompts, shown in Table 10 and 11, are used consistently in all experimental evaluations of WebQSP and CWQ.
@@ Appendix E Additional Results
E.1 Influence of Reasoning Paths Number
@@
E.2 Reasoning Process Quality and Visualizations (Additional Results)

### v2

@@ 4.1 Experimental Setups
We conduct comprehensive experiments to evaluate the effectiveness and generalization ability of KG-TRACES across both general-domain and domain-specialized (medicine) reasoning tasks. We evaluate the general reasoning ability of KG-TRACES on two open-domain multi-hop KGQA benchmarks: WebQuestionsSP (WebQSP) [41] and Complex WebQuestions (CWQ) [29]. To evaluate cross-domain generalization, we additionally use GenMedGPT-5k [15], a medical QA benchmark constructed from ChatGPT-patient interactions and supported by a curated medical KG of symptoms, diseases, drugs, and treatments [34]. To validate the quality of our generated data without an explicit filtering stage, we conducted a post-hoc evaluation using a panel of powerful LLMs (GPT-4o, Claude-4-sonnet, and Gemini-2.5-Pro). Across 2,000 samples, our dataset achieved high average scores of 8.1/10 on WebQSP and 7.4/10 on CWQ, confirming its high fidelity. This high quality was further corroborated by a rigorous manual verification on 150 samples (50 from WebQSP, 100 from CWQ), where human evaluation confirmed that 86% and 83% of the generated reasoning processes, respectively, were logically sound and factually correct. The details of the datasets are provided in Appendix A.
@@ A.1 General Reasoning Dataset
To demonstrate the high quality of the reasoning processes generated by Qwen-72B-Instruct, we performed a random sample post-hoc validation. This audit serves to quantify the reliability and coherence of our automatically constructed dataset, justifying its direct use for model training without an intermediate filtering step.
Our validation methodology involved a committee of powerful, closed-source LLMs (GPT-4o, Claude-4-sonnet, and Gemini-2.5-Pro) acting as evaluators. We randomly sampled 500 instances from the WebQSP subset and 1,500 from the CWQ subset. Each instance, comprising the question, ground truth answer, and the generated reasoning process, was scored by the evaluators on a scale of 0-10.
The prompt used for this validation is presented in Table 4. It was designed to be simple and direct, focusing on the core correctness and quality of the reasoning.
| Score Prompt (Reasoning Process Quality) |
| Given the following question, ground truth answer, and a reasoning process, please evaluate the correctness and quality of the reasoning on a scale of 0 to 10. Provide only the score as an integer. Question: {question} |
| Ground Truth Answer: {ground truth answer} |
| Reasoning process: {llm generated reasoning process} |
@@ B.2 Inference Prompts
These inference prompts, shown in Table 11 and 12, are used consistently in all experimental evaluations of WebQSP and CWQ.
@@ Appendix E Additional Results
E.1 Granular Performance Analysis
In this section, we provide a more granular analysis of our model’s performance on the WebQSP and CWQ test sets, broken down by the complexity of the questions. This analysis addresses the model’s scalability on multi-hop reasoning and its robustness to questions with varying numbers of answers.
E.1.1 Analysis by Hop Depth
As shown in Table 14, our model demonstrates strong performance across different hop depths. Notably, on the most complex >=3 hop questions in CWQ, KG-TRACES significantly outperforms prior SOTA methods, validating its effectiveness on deeper reasoning chains.
| Methods | WebQSP | CWQ | ||||
| 1 hop | 2 hop | hop | 1 hop | 2 hop | hop | |
| RoG | 77.03 | 64.86 | - | 62.88 | 58.46 | 37.82 |
| KG-TRACES (ours) | 80.53 | 67.23 | - | 64.26 | 58.73 | 45.37 |
E.1.2 Analysis by Answer Number
Table 15 shows the model’s performance on questions with different numbers of ground-truth answers. This demonstrates the model’s robustness in handling both single-answer and multi-answer scenarios.
| Methods | WebQSP | CWQ | ||||||
| #Ans=1 | 2#Ans4 | 5#Ans9 | #Ans10 | #Ans=1 | 2#Ans4 | 5#Ans9 | #Ans10 | |
| RoG | 67.89 | 79.39 | 75.04 | 58.33 | 56.90 | 53.73 | 58.36 | 43.62 |
| KG-TRACES (ours) | 74.04 | 84.65 | 76.87 | 57.87 | 60.68 | 53.07 | 60.07 | 46.56 |
E.2 Influence of Reasoning Paths Number
@@
E.3 Reasoning Process Quality and Visualizations (Additional Results)